OpenStudy (anonymous):

Find the simplified form of the expression. Write your answer in scientific notation. (3x10^6) (8x10^-4)

OpenStudy (ybarrap):

(3x10^6) (8x10^-4) =(3*8) * (10^6 * 10^4) =24 * 10^(6 - 4) =24 * 10^(2) =2.4 * 10^(3) Make sense?